New Jersey Valve Company manufactured 7,000 units during January of a control valve used by milk processors in its Camden plant. Records indicated the following:

Direct labor 36,100 hr. at $14.70 per hr.
Direct material purchased 28,000 lb. at $2.20 per lb.
Direct material used 26,000 lb.

The control valve has the following standard prime costs:

     
Direct material 4 lb. at $2.10 per lb. $ 8.40
Direct labor 5 hr. at $15.10 per hr. 75.50
Standard prime cost per unit $ 83.90

Required:

  1. Prepare a schedule of standard production costs for January, based on actual production of 7,000 units.
  2. For the month of January, compute the following variances.

Prepare a schedule of standard production costs for January, based on actual production of 7,000 units.

NEW JERSEY VALVE COMPANY: CAMDEN PLANT
Schedule of Standard Production Costs
Based on 7,000 Units
For the Month of January
Standard Costs
Direct material
Direct labor
Total standard production costs

For the month of January, compute the following variances. (Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).)

a. Direct-material price variance   
b. Direct-material quantity variance
c. Direct-material purchase price variance
d. Direct-labor rate variance
e. Direct-labor efficiency variance
Answer #1
Schedule of standard production costs - New Jersey valve Company
Particulars cost per unit Total cost
Direct materials (7000X8.40) 8.4 58,800.00
Direct labor (7000X 75.5) 75.5 528,500.00
Total standard production costs 587,300.00
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($2.10 - $2.20 ) × 28000 pound = $2800 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((4Pound   X 7000 Unit)-26000 Pound )X $2.10 = $4200 F
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($15.10/hour - $14.70/hour) × 36100 Hours= $14440 F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (5Hour × 7000Unit - 36100) × $15.10 = $16610 U F